OpenStudy (anonymous):

Multiply and simplify. Show work/steps (x-3)(x-4)

OpenStudy (anonymous):

( x + a ) ( x + b ) = x^2 + ( a + b ) x + ab ( x + 3 ) ( x - 4 ) here a = 3 and b= -4 then ( x + 3 ) ( x - 4 ) = x^2 + ( 3 - 4 ) x + 3 * ( - 4 ) .........................= x^2 - x - 12

OpenStudy (anonymous):

x(x-4)-3(x-4) EXPAND :)

OpenStudy (anonymous):

x + a ) ( x + b ) = x^2 + ( a + b ) x + ab ( x + 3 ) ( x - 4 ) here a = 3 and b= -4 then ( x + 3 ) ( x - 4 ) = x^2 + ( 3 - 4 ) x + 3 * ( - 4 ) ( x + 3 ) ( x - 4 )= x^2 - x - 12 Now award a metal or best response

OpenStudy (anonymous):

I see! Thank you! I see where I went wrong.